Find the solution of each of the following initial-value problems.

(1)=60x (55%*)+*(== 3)-**
Answer #1

9) X= (-3)* + Comm], 409-[!] where *= ( x ) and X - [3 ] -> (3]-[ -=][:]+[ace v ] -C] ( 37 600) - x1 = 2x-24 . X(0), yco)=0L{x} = ( (see) -2 et (S-2) (342) 48 St2) -20 TS L{x} = (2+2)-20 <= y s²4 + 8 stem 2xy = $ +2-se 23 - Stey & Sell Sell -AS ya= 1 . . . . ? uz 109.13–1986324 ya ile 3 (19_222* 00sinal) yet)= L = TVE13 - ILe C ycke sinalom By en verse Laplace second